Find the probability of exactly one 5 when a die is rolled 3 times 2) Find the probability of getting 3 heads when 8 coins are tossed. 3) A bag contains 4 red and 2 green balls. A ball is drawn and replaced 4 times. What is the probability of getting exactly 3 red balls and 1 green ball.


1
Expert's answer
2021-06-01T13:06:59-0400

Find the probability of exactly one 5 when a die is rolled 3 times 

"P(1)" Probability of exactly 1 successes

"\text{trials} = 3"

"p = \frac{1}{6}=0.167" and "X = 1"

into a binomial probability distribution function (PDF). If doing this by hand,

apply the binomial probability formula:

"P(X) = \binom{n}{X} \cdot p^X \cdot (1-p)^{n-X}"

The binomial coefficient,"\binom{n}{X}" is defined by

"\binom{n}{X} = \frac{n!}{X!(n-X)!}"


The full binomial probability formula with the binomial coefficient is

"P(X) = \frac{n!}{X!(n-X)!} \cdot p^X \cdot (1-p)^{n-X}"


where n is the number of trials, p is the probability of success on a single trial, and X

is the number of successes. Substituting in values for this problem, 


n=3

p=0.167

X=1

"P(1) = \frac{3!}{1!(3-1)!} \cdot 0.167^1 \cdot (1-0.167)^{3-1}"

Evaluating the expression, we have

P(1)=0.347638389




2) Find the probability of getting 3 heads when 8 coins are tossed.  

"P(3)" Probability of exactly 3 successes

"\text{trials} = 8"

"p = \frac{3}{8}=0.375" and "X = 3"

into a binomial probability distribution function (PDF). If doing this by hand,

apply the binomial probability formula:

"P(X) = \binom{n}{X} \cdot p^X \cdot (1-p)^{n-X}"

The binomial coefficient,"\binom{n}{X}" is defined by

"\binom{n}{X} = \frac{n!}{X!(n-X)!}"


The full binomial probability formula with the binomial coefficient is

"P(X) = \frac{n!}{X!(n-X)!} \cdot p^X \cdot (1-p)^{n-X}"


where n is the number of trials, p is the probability of success on a single trial, and X

is the number of successes. Substituting in values for this problem, 


n=8

p=0.375

X=3

"P(3) = \frac{8!}{3!(8-3)!} \cdot 0.375^3 \cdot (1-0.375)^{8-3}"

Evaluating the expression, we have

P(3)=0.28163194656372


3) A bag contains 4 red and 2 green balls. A ball is drawn and replaced 4 times. What is the probability of getting exactly 3 red balls and 1 green ball.

we find probability for exact 1 green ball then automatically remaining balls are red.

"P(1)" Probability of exactly 1 successes

"\text{trials} = 4"

"p = \frac{2}{6}=0.333" and "X = 1"

into a binomial probability distribution function (PDF). If doing this by hand,

apply the binomial probability formula:

"P(X) = \binom{n}{X} \cdot p^X \cdot (1-p)^{n-X}"

The binomial coefficient,"\binom{n}{X}" is defined by

"\binom{n}{X} = \frac{n!}{X!(n-X)!}"


The full binomial probability formula with the binomial coefficient is

"P(X) = \frac{n!}{X!(n-X)!} \cdot p^X \cdot (1-p)^{n-X}"


where n is the number of trials, p is the probability of success on a single trial, and X

is the number of successes. Substituting in values for this problem, 


n=4

p=0.333

X=1

"P(1) = \frac{4!}{1!(4-1)!} \cdot 0.333^1 \cdot (1-0.333)^{4-1}"

Evaluating the expression, we have

P(1)=0.395258962716